OpenStudy (anonymous):

solve the indicated set: (3x+5)(2x-8)=0

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

factoring?

OpenStudy (skullpatrol):

You are on the right track...

OpenStudy (anonymous):

6x^2 -14x+40=0

OpenStudy (anonymous):

-40

OpenStudy (skullpatrol):

...but you already have two factors: one is (3x+5) and the other is (2x-8)

OpenStudy (anonymous):

so do you make both of them equal to 0?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

will the answers be rational reals?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

x=4, -5/3

OpenStudy (skullpatrol):

Yes, either x=4 or x=-5/3 will make (3x+5)(2x-8)=0 a true statement, right?

OpenStudy (skullpatrol):

Can you tell me why?

OpenStudy (anonymous):

Is it just an accident that the opposite of the numbers divided is the answer? ex. (2x-8) -8 and 2

OpenStudy (skullpatrol):

x=4 or x=-5/3 are the solutions to the equation: $$\Huge (3x+5)(2x-8)=0 $$

OpenStudy (skullpatrol):

Let x = 4, substitute that value of x into the equation, what do you get?

OpenStudy (anonymous):

0

OpenStudy (skullpatrol):

Yes, you get 0 on the left hand side because 0 times any number is 0, and because there is a 0 on the right hand side you get the true statement 0 = 0, right?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

The same applies for the other solution x = -5/3

OpenStudy (skullpatrol):

Thus the solution set is {4, -5/3}

OpenStudy (anonymous):

ok thank you!

OpenStudy (skullpatrol):

Thanks for trying to learn.